# Application of Optimal Control to Time-Resolution Protocol for Quantum Sensing

### v1

@@ III.2 Short interrogation time $\tau<T^{*}$ I: detune protocol in RWA
$e^{-iH_{\text{RWA}}\tau}=\cos(\frac{\tilde{\Omega}\tau}{2})\,{\color[rgb]{0,0,0}\hat{e}_{2\times 2}}-i\sin(\frac{\tilde{\Omega}\tau}{2})\bigg{[}\frac{\Delta}{\tilde{\Omega}}\hat{\sigma}_{z}+\frac{u_{\text{max}}}{\tilde{\Omega}}\big{(}\hat{\sigma}_{x}\cos\theta+\hat{\sigma}_{y}\sin\theta\big{)}\bigg{]}$ (23)
@@
To provide some intuition, in Fig. 3 we visualize how the detune protocol enhances sensitivity by comparing trajectories of two detune protocols on the Bloch sphere where the sensor qubit state in rotating frame is parametrized by $|\tilde{\psi}_{0}\rangle=\big{[}\cos\frac{\theta}{2},\sin\frac{\theta}{2}e^{i\phi}\big{]}^{T}$. Within RWA, the detune protocol has a time-independent Hamiltonian and each trajectory is associated with a rotating axis of the Bloch sphere. Using Eq. (22), given a detune $\Delta\omega$ and a to-be-measured field $\delta\omega$, the rotating axis is pointing along $(0,u_{\text{max}},\delta\omega-\Delta\omega)$. The square of projection $\cos^{2}(\frac{\theta}{2})$ at the terminal point of the trajectory is the measurement signal. As a concrete example we consider $u_{\text{max}}=0.2$ and $\tau=\frac{t_{\text{QSL}}}{2}$. For the zero detune $\Delta\omega=0$, the terminal points of two trajectories with $\delta\omega=0$ and 0.1 [blue curves in Fig. 3(a)] are both very close to the equator of Bloch sphere. Their projections on $|0\rangle$ are close [blue curves of Fig. 3(b)], indicating a low sensitivity for the field strength of $\delta\omega=0.1$. When using the optimal detune where $\Delta\omega\sim 0.326$, the terminal point of $\delta\omega=0.1$ trajectory is close to equator [dash red curve in Fig. 3(a)] but that of $\delta\omega=0$ trajectory is not [solid red curve in Fig. 3(a)]. Their projections on $|0\rangle$ are more distinct [red curves of Fig. 3(b)], indicating an enhanced sensitivity for the field strength of $\delta\omega=0.1$.
@@ III.3 Short interrogation time $\tau<T^{*}$ II: full calculation
We now use full calculation to obtain the best detune and optimal protocols in the $\tau<t_{\text{QSL}}$ regime. With the full calculation, we re-parametrize the detune protocol [which is completely equivalent to Eq. (21)] as
@@ III.4 Long interrogation time $\tau>T^{*}$
For completeness we consider the $\tau>T^{*}$ regime which is less relevant to high time resolution. In this regime the behavior is qualitatively similar to the celebrated Ramsey protocol Ramsey (1950). In Fig. 5 we show the results for $u_{\text{max}}=0.1$, 0.5 with the interrogation time $\tau=1.2\,t_{\text{QSL}}$.
The optimal control can be divided into three stages. In the first stage the optimal control brings the sensor qubit to the equal-superposition of two natural eigenstates (equator on the Bloch sphere) using BB waveform in a very short time; this corresponds to the first $\frac{\pi}{2}$ pulse in Ramsey protocol and is demonstrated by the reduction of $\langle\sigma_{z}\rangle(t)$ from 1 to (near) 0 shown in Fig. 5(a), (c). The second stage is an S control which corresponds to the free evolution in the Ramsey protocol. One notes that in the optimal protocol the S control is not necessarily zero; particularly it can be significantly different from zero for large $u_{\text{max}}$ as shown in Fig. 5(c). The final stage is again a BB control that maximizes the projection difference with and without $\delta\omega$. Overall for small $\frac{u_{\text{max}}}{\omega_{0}}$, the optimal protocol does behave like the Ramsey protocol; the main difference is that the initial and final $\frac{\pi}{2}$ pulses in the Ramsey sequence are replaced by the faster BB protocols Lin et al. (2025).
@@ IV.1 Smoothness and deviation from optimal performance
In this section, we address two practical issues related to short-$\tau$ time-resolution protocols. The short-$\tau$ optimal protocol is purely BB which contains discontinuities at switching times, and one can smooth the protocol by introducing a running cost that imposes smoothness. The degree of smoothness can be quantified by Lin et al. (2025)
@@ IV.2 Application to superconducting qubit
Figure 7: Potential application in calibrating the Z pulse distortion in superconducting qubits. The distortion can have time constants ranging from tens of nanoseconds to microseconds Rol et al. (2020); Hellings et al. (2025); Li et al. (2025). The desired waveform is a step pulse (known) shown in blue and the actual waveform experienced by the qubit is a distorted pulse (unknown) shown in orange. Both are shown in arbitrary unit. To calibrate the distortion, we first need to reconstruct the distorted waveform. We apply the short-$\tau$ time-resolution protocol centered at a series of times $t$ following the onset of the reference waveform, repeating each measurement many times to build sufficient statistics. For $u_{\text{max}}=0.4$ GHz, the time resolution can be made smaller than $\tau\sim t_{\text{QSL}}=\frac{\pi}{u_{\text{max}}}=8$ ns. In the plot $\tau\sim 20$ ns is used.
@@
As stated at the end of Section II.1, employing the time-resolution protocol requires a system capable of repeatedly generating the identical, yet unknown time-domain signal to be measured. Generating identical time-domain signals implies the ability of preparing the system of interest in exactly the same (but not necessarily known) initial state, which is the primary factor limiting its practical application. For example if the goal is to measure the magnetic domain wall propagation using the time-resolution protocol Herb et al. (2025), one in principle needs to repeatedly prepare the system with a magnetic domain structure with nanometer-scale precision, which can be challenging. With the hardware constraints in mind, in the following we outline a potentially impactful application.
@@
In the context of superconducting qubits, baseband flux pulses are widely used to implement quantum gates such as the physical $Z$ gate, the iSWAP gate, and the controlled-phase (CPHASE) gate Kjaergaard et al. (2022); Negîrneac et al. (2021); Jin et al. (2015); Barends et al. (2014); Krantz et al. (2019); Ding et al. (2025); Chen et al. (2022); Collodo et al. (2020); Chu and Yan (2021); Kelly et al. (2014); Sung et al. (2021); Zhao et al. (2023). A long-standing challenge in these flux-based control schemes is pulse distortion – the deviation between the intended flux waveform and the actual waveform experienced by the qubit, arising from issues such as filtering effects and impedance mismatches in the control lines. To address this, we propose using the time-resolution protocol (preferably the detune variant) to characterize and compensate for these distortions. The procedure is illustrated in Fig. 7. In the example shown, the desired waveform is a known step pulse (blue, dashed curve), while the actual waveform experienced by the qubit is an unknown, distorted pulse (orange, solid curve). Good distortion calibration necessitates an accurate reconstruction of the distorted waveform. We apply the short-$\tau$ time-resolution protocol centered at a series of times $t$ following the onset of the reference waveform, repeating each measurement many times to build sufficient statistics. Moreover since the commonly used dispersive readout of superconducting qubits belongs to the Quantum nondemolition (QND) measurement Blais et al. (2004, 2021); Lupascu et al. (2007), i.e., the qubit remains in one of natural eigenstates following each short-$\tau$ measurement, we can sequentially probe the same distorted waveform at different values of $t$, assuming the protocol does not significantly alter the profile of the distorted pulse. A complete calibration procedure using sequential measurements could substantially reduce the time required to collect sufficient data, thereby accelerating the overall calibration process.
@@ V Conclusion
We apply OCT to analyze the optimal performance of time resolution in quantum sensing. The performance of time resolution is quantified by the sensitivity with respect to the external field to be measured in a short interrogation time $\tau$. For this task, an end-to-end description of the sensor state over the entire measurement process is essential as any operation contributes to the interrogation time. In particular, the initial state and the basis of projection measurement have to be in the natural eigenstates of the sensor qubit. One theoretical/numerical complication arising from time resolution when applying OCT is that the terminal cost involves the derivative of the quantum state, and we cope with this additional dependence by augmenting the dynamics. This procedure not only facilitates the numerics but keeps the control system time-invariant and control-affine so that the optimality conditions from OCT remains valid.
Similar to the reference YX protocol proposed in Ref. Herb and Degen (2024), the behavior of optimal protocols exhibits a critical interrogation time $T^{*}$: when $\tau<T^{*}$ the optimal protocol is purely bang-bang; when $\tau>T^{*}$ the optimal protocol involves a singular control the interrogation. The latter corresponds to the free evolution in the Ramsey protocol. Given the same amplitude constraint $u_{\text{max}}$, the optimal control achieves a higher sensitivity over the reference protocol. The metrological gain is more substantial in the regime of short interrogation time, the regime relevant to high time resolution.
@@ Appendix A Kernel function construction
In this Appendix we provide a straightforward way to construct the kernel $K$ in Eq. (8). Without loss of generality we take $t=0$ to simplify the notation. $K(t^{\prime})$ vanishes when $|t^{\prime}|>\frac{\tau}{2}$ because during these periods the sensor qubit is in one of natural eigenstates (i.e., eigenstates of $\hat{\sigma}_{z}$) that does not respond to $\delta\omega\,\hat{\sigma}_{z}$. When $\delta\omega(t^{\prime})=\delta\omega$ is a constant over $|t^{\prime}|<\frac{\tau}{2}$, the integrated area $\int\text{d}t^{\prime}K(t^{\prime})=\eta$ .
@@
given in Ref. Herb and Degen (2024). Results for $\tau=0.5\,t_{\text{QSL}}$ for YX and detune protocols using RWA are given in Fig. 8. It is seen that the kernel of the detune protocol is smoother. The integrated areas of both protocols are comparable, indicating they have similar sensitivity values.

### v2

@@ III.2 Short interrogation time $\tau<T^{*}$ I: detune protocol in RWA
$e^{-iH_{\text{RWA}}\tau}=\cos(\frac{\tilde{\Omega}\tau}{2})\,\hat{e}_{2\times 2}-i\sin(\frac{\tilde{\Omega}\tau}{2})\bigg[\frac{\Delta}{\tilde{\Omega}}\hat{\sigma}_{z}+\frac{u_{\text{max}}}{\tilde{\Omega}}\big(\hat{\sigma}_{x}\cos\theta+\hat{\sigma}_{y}\sin\theta\big)\bigg]$ (23)
@@
To provide some intuition, in Fig. 3 we visualize how the detune protocol enhances sensitivity by comparing trajectories of two detune protocols on the Bloch sphere where the sensor state in rotating frame is parametrized by $|\tilde{\psi}_{0}\rangle=\big[\cos\frac{\theta}{2},\sin\frac{\theta}{2}e^{i\phi}\big]^{T}$. Within RWA, the detune protocol has a time-independent Hamiltonian and each trajectory is associated with a rotating axis of the Bloch sphere. Using Eq. (22), given a detune $\Delta\omega$ and a to-be-measured field $\delta\omega$, the rotating axis is pointing along $(0,u_{\text{max}},\delta\omega-\Delta\omega)$. The square of projection $\cos^{2}(\frac{\theta}{2})$ at the terminal point of the trajectory is the measurement signal. As a concrete example we consider $u_{\text{max}}=0.2$ and $\tau=\frac{t_{\text{QSL}}}{2}$. For the zero detune $\Delta\omega=0$, the terminal points of two trajectories with $\delta\omega=0$ and 0.1 [blue curves in Fig. 3(a)] are both very close to the equator of Bloch sphere. Their projections on $|0\rangle$ are close [blue curves of Fig. 3(b)], indicating a low sensitivity for the field strength of $\delta\omega=0.1$. When using the optimal detune where $\Delta\omega\sim 0.326$, the terminal point of $\delta\omega=0.1$ trajectory is close to equator [dash red curve in Fig. 3(a)] but that of $\delta\omega=0$ trajectory is not [solid red curve in Fig. 3(a)]. Their projections on $|0\rangle$ are more distinct [red curves of Fig. 3(b)], indicating an enhanced sensitivity for the field strength of $\delta\omega=0.1$.
@@ III.3 Short interrogation time $\tau<T^{*}$ II: full calculation
We now use full calculation to obtain the best detune and optimal protocols in the $\tau<t_{\text{QSL}}$ regime. With the full calculation, we re-parametrize the detune protocol [completely equivalent to Eq. (21)] as
@@ III.4 Long interrogation time $\tau>T^{*}$
For completeness we consider the $\tau>T^{*}$ regime which is less relevant to high time resolution. In this regime the behavior is qualitatively similar to the celebrated Ramsey protocol PhysRev.78.695. In Fig. 5 we show the results for $u_{\text{max}}=0.1$, 0.5 with the interrogation time $\tau=1.2\,t_{\text{QSL}}$.
The optimal control can be divided into three stages. In the first stage the optimal control brings the quantum sensor to the equal-superposition of two natural eigenstates (equator on the Bloch sphere) using BB waveform in a very short time; this corresponds to the first $\frac{\pi}{2}$ pulse in Ramsey protocol and is demonstrated by the reduction of $\langle\sigma_{z}\rangle(t)$ from 1 to (near) 0 shown in Fig. 5(a), (c). The second stage is an S control which corresponds to the free evolution in the Ramsey protocol. One notes that in the optimal protocol the S control is not necessarily zero; particularly it can be significantly different from zero for large $u_{\text{max}}$ as shown in Fig. 5(c). The final stage is again a BB control that maximizes the projection difference with and without $\delta\omega$. Overall for small $\frac{u_{\text{max}}}{\omega_{0}}$, the optimal protocol does behave like the Ramsey protocol; the main difference is that the initial and final $\frac{\pi}{2}$ pulses in the Ramsey sequence are replaced by the faster BB protocols PhysRevA.111.042602.
@@ IV.1 Smoothness and deviation from optimal performance
In this section, we address two practical issues related to short-$\tau$ time-resolution protocols. The short-$\tau$ optimal protocol is purely BB which contains discontinuities at switching times, and one can smooth the protocol by introducing a running cost that imposes smoothness. The degree of smoothness can be quantified by PhysRevA.111.042602
@@ IV.2 Application to superconducting qubit
Figure 7: Potential application in calibrating the Z pulse distortion in superconducting qubits. The distortion can have time constants ranging from tens of nanoseconds to microseconds 10.1063/1.5133894; hellings2025calibratingmagneticfluxcontrol; PhysRevApplied.23.024059. The desired waveform is a step pulse (known) shown in blue and the actual waveform experienced by the qubit is a distorted pulse (unknown) shown in orange. Both are shown in arbitrary unit. To calibrate the distortion, we first need to reconstruct the distorted waveform. We apply the short-$\tau$ time-resolution protocol centered at a series of times $t$ following the onset of the reference waveform, repeating each measurement many times to build sufficient statistics. For $u_{\text{max}}=0.4$ GHz, the time resolution can be made smaller than $\tau\sim t_{\text{QSL}}=\frac{\pi}{u_{\text{max}}}=8$ ns. In the plot $\tau\sim 20$ ns is used.
@@
As stated at the end of Section II.1, employing the time-resolution protocol requires a system capable of repeatedly generating the identical, yet unknown time-domain signal to be measured. Generating identical time-domain signals implies the ability of preparing the system of interest in exactly the same (but not necessarily known) initial state, which is the primary factor limiting its practical application. For example if the goal is to measure the magnetic domain wall propagation using the time-resolution protocol Herb_2025, one in principle needs to repeatedly prepare the system with a magnetic domain structure with nanometer-scale precision, which can be challenging. With the hardware constraints in mind, in the following we outline a potentially impactful application.
@@
In the context of superconducting qubits, baseband flux pulses are widely used to implement quantum gates such as the physical $Z$ gate, the iSWAP gate, and the controlled-phase (CPHASE) gate PhysRevX.12.011005; negirneac_high_fidelity_2021; PhysRevApplied.3.034004; barends_superconducting_2014; Krantz_2019; yskp-mfcr; PhysRevApplied.18.034027; Collodo2020; chu_coupler_assisted_2021; kelly_prl_2014; PhysRevX.11.021058; PhysRevApplied.19.054050. A long-standing challenge in these flux-based control schemes is pulse distortion – the deviation between the intended flux waveform and the actual waveform experienced by the qubit, arising from issues such as filtering effects and impedance mismatches in the control lines. To address this, we propose using the time-resolution protocol (preferably the detune variant) to characterize and compensate for these distortions. The procedure is illustrated in Fig. 7. In the example shown, the desired waveform is a known step pulse (blue, dashed curve), while the actual waveform experienced by the qubit is an unknown, distorted pulse (orange, solid curve). Good distortion calibration necessitates an accurate reconstruction of the distorted waveform. We apply the short-$\tau$ time-resolution protocol centered at a series of times $t$ following the onset of the reference waveform, repeating each measurement many times to build sufficient statistics. Moreover since the commonly used dispersive readout of superconducting qubits belongs to the Quantum nondemolition (QND) measurement PhysRevA.69.062320; RevModPhys.93.025005; Lupascu-2007, i.e., the qubit remains in one of natural eigenstates following each short-$\tau$ measurement, we can sequentially probe the same distorted waveform at different values of $t$, assuming the protocol does not significantly alter the profile of the distorted pulse. A complete calibration procedure using sequential measurements could substantially reduce the time required to collect sufficient data, thereby accelerating the overall calibration process.
@@ V Conclusion
We apply OCT to analyze the optimal performance of time resolution in quantum sensing. The performance of time resolution is quantified by the sensitivity with respect to the external field to be measured in a short interrogation time $\tau$. For this task, an end-to-end description of the sensor state over the entire measurement process is essential as any operation contributes to the interrogation time. In particular, the initial state and the basis of projection measurement have to be in the natural eigenstates of the quantum sensor. One theoretical/numerical complication arising from time resolution when applying OCT is that the terminal cost involves the derivative of the quantum state, and we cope with this additional dependence by augmenting the dynamics. This procedure not only facilitates the numerics but keeps the control system time-invariant and control-affine so that the optimality conditions from OCT remains valid.
Similar to the reference YX protocol proposed in Ref. PhysRevLett.133.210802, the behavior of optimal protocols exhibits a critical interrogation time $T^{*}$: when $\tau<T^{*}$ the optimal protocol is purely bang-bang; when $\tau>T^{*}$ the optimal protocol involves a singular control the interrogation. The latter corresponds to the free evolution in the Ramsey protocol. Given the same amplitude constraint $u_{\text{max}}$, the optimal control achieves a higher sensitivity over the reference protocol. The metrological gain is more substantial in the regime of short interrogation time, the regime relevant to high time resolution.
@@ Appendix A Kernel function construction
In this Appendix we provide a straightforward way to construct the kernel $K$ in Eq. (8). Without loss of generality we take $t=0$ to simplify the notation. $K(t^{\prime})$ vanishes when $|t^{\prime}|>\frac{\tau}{2}$ because during these periods the quanutm sensor is in one of natural eigenstates (i.e., eigenstates of $\hat{\sigma}_{z}$) that does not respond to $\delta\omega\,\hat{\sigma}_{z}$. When $\delta\omega(t^{\prime})=\delta\omega$ is a constant over $|t^{\prime}|<\frac{\tau}{2}$, the integrated area $\int\text{d}t^{\prime}K(t^{\prime})=\eta$ .
@@
given in Ref. PhysRevLett.133.210802. Results for $\tau=0.5\,t_{\text{QSL}}$ for YX and detune protocols using RWA are given in Fig. 8. It is seen that the kernel of the detune protocol is smoother. The integrated areas of both protocols are comparable, indicating they have similar sensitivity values.